Find the zeros of the function.

f(x) = 6x3 + 18x2 + 12x

help me please
1
Expert's answer
2014-06-06T04:20:42-0400
We can write the function as f(x)=x(6x^2+18x+12).
From this we can see that one zero is x=0. To find others we need to solve the next quadratic equations x^2+3x+2=0.

The zeros are: x_1=(-3+(9-8)^(1/2))/2=(-3+1)/2=-1 and x_2=(-3-1)/2=-2.

Answer. The zeros are: 0, -1 and -2.